# Combined Pre-Supernova Alert System with KamLAND and Super-Kamiokande

### v2

@@ Abstract
Preceding a core-collapse supernova, various processes produce an increasing amount of neutrinos of all flavors characterized by mounting energies from the interior of massive stars. Among them, the electron antineutrinos are potentially detectable by terrestrial neutrino experiments such as KamLAND and Super-Kamiokande via inverse beta decay interactions. Once these pre-supernova neutrinos are observed, an early warning of the upcoming core-collapse supernova can be provided. In light of this, KamLAND and Super-Kamiokande have been monitoring pre-supernova neutrinos since 2015 and 2021, respectively. Recently, we performed a joint study between KamLAND and Super-Kamiokande on pre-supernova neutrino detection. A pre-supernova alert system combining the KamLAND detector and the Super-Kamiokande detector is developed and put into operation, which can provide a supernova alert to the astrophysics community. Fully leveraging the complementary properties of these two detectors, the combined alert is expected to resolve a pre-supernova neutrino signal from a 15 M⊙ star within 510 pc of the Earth, at a significance level corresponding to a false alarm rate of no more than 1 per century. For a Betelgeuse-like model with optimistic parameters, it can provide early warnings up to 12 hours in advance.
@@ 1 Introduction
Neutrinos emitted by a supernova during the first 10 seconds carry unique information about the physics of supernovae, which hold immense significance in the realm of astrophysics. The first observed supernova neutrinos were from SN1987A in the Large Magellanic Cloud, 50 kpc away from Earth (Hirata et al., 1987; Bionta et al., 1987; Alekseev et al., 1987). Since then, various neutrino experiments, such as Borexino (Alimonti et al., 2009), IceCube (Köpke, 2018), KamLAND (Abe et al., 2022b), LVD (Agafonova et al., 2015), NOvA (Acero et al., 2020), SNO+ (Andringa et al., 2016), and Super-Kamiokande (SK) (Abe et al., 2016), equipped with advanced technology and improved capabilities, continued the quest to detect supernova neutrino bursts. Furthermore, a number of next-generation neutrino detectors sensitive to galactic supernova neutrinos are under construction, including DUNE (Abi et al., 2021), Hyper-Kamiokande (Abe et al., 2018), JUNO (Abusleme et al., 2023), and KM3NeT (Aiello et al., 2021). To catch such a fleeting event, it is desirable that we be alerted well before the explosion, so that astronomers and particle physicists can get ready to observe the supernova neutrinos and possible gravitational waves as soon as the explosion happens.
The evolution of a star, whose initial mass is greater than 8 solar masses (M⊙), to its final stages prior to the core-collapse supernova (CCSN) is characterized by the nuclear burning in its core due to the high temperature and density (Woosley et al., 2002). The change of the chemical composition of a star, forming concentric shells of heavier elements along its volume, is the result of nuclear fusion of heavier elements in the core. Stars at this stage are called pre-supernova (pre-SN) stars. The main cooling mechanism of a pre-SN star is through photon and neutrino emissions. Neutrinos are mainly produced by pair annihilation , producing all flavours of neutrino and antineutrino pairs (Odrzywolek et al., 2004a). As the star approaches core collapse, the nuclear beta decay begins to dominate. Nuclear processes, such as beta decay, will eventually contribute more to the neutrino emission than thermal processes as the star is approaching core collapse (Patton et al., 2017).
@@
The energy of pre-SN neutrinos is of sub-MeV scale or MeV scale. We focus on inverse beta decay (IBD) , which has a relatively large cross section in liquid scintillator (LS) detectors (e.g. KamLAND (Suzuki, 2014)) and water Cherenkov detectors (e.g. SK (Fukuda et al., 2003)) in the energy range of pre-SN neutrinos. Compared to LS detectors, water Cherenkov detectors are less sensitive to low-energy neutrinos because the higher energy threshold limits the detection of neutron capture signals. However, starting in 2020, the SK detector was loaded with gadolinium (Gd) to improve the neutron detection efficiency (Abe et al., 2022a). Since 2015, KamLAND has been monitoring pre-SN neutrinos and able to provide pre-SN alert to the astrophysics community (Asakura et al., 2016). Later in 2021, SK has also put a pre-SN alert system online (Machado et al., 2022). Besides IBD in LS detectors and water Cherenkov detectors, distinct detection methods in other detectors could be utilized for pre-SN neutrino detection as well. For example, coherent neutrino-nucleus scattering in future large scale dark matter direct detection experiments, is a method complementary to IBD, because it can detect all flavors of neutrinos (Raj et al., 2020).
In this article, we introduce a joint study between KamLAND and SK on pre-SN neutrino detection. This combination aims at extending the reach to potential CCSN progenitors further away and the warning time of pre-SN alerts. Additionally, we present new sensitivity results individually for KamLAND and SK. Compared to the previous study (Asakura et al., 2016), KamLAND has now taken newer pre-SN neutrino models into consideration. The sensitivity of SK to pre-SN neutrinos with Gd by mass is presented in (Machado et al., 2022). Since 2022, the Gd concentration in SK has increased to 0.03 by mass, further enhancing its capability to identify low-energy electron antineutrinos () via IBD. Therefore, both experiments have reoptimized their analysis strategies according to the changes.
@@ 2 Pre-supernova neutrino model
To estimate the expected signal from pre-SN in SK and KamLAND, two models for pre-SN neutrino emission during the evolution of massive stars were used: (Odrzywolek & Heger, 2010; Odrzywolek et al., 2004b) and (Patton et al., 2017). Both models provide data sets for the calculation of emission during the pre-SN stage. (Odrzywolek et al., 2004b) provides the data sets for stars with 15 M⊙ and 25 M⊙ and (Patton et al., 2017) for 15 M⊙ and 30 M⊙ .
@@
The model from (Odrzywolek & Heger, 2010) assumes that the entire neutrino flux comes from pair annihilation. For the nuclear isotopic composition of the star, the model assumes a nuclear statistical equilibrium (NSE), which is a treatment only dependent on the temperature, density, and electron fraction, making it a simple flux estimated by only post-processing an already existing stellar model. The model from (Patton et al., 2017) includes a more complete evaluation of the neutrino flux from the pre-SN star, including contributions not only from pair annihilation, but also from plasmon decay, photoneutrino process, beta decay, and electron capture. By using the star evolution code MESA (Modules for Experiments in Stellar Astrophysics) (Paxton et al., 2011), this model couples the isotopic evolution to the stellar evolution, giving a more robust estimation of the neutrino fluxes from nuclear weak processes.
@@
To calculate the expected signal from the considered models, an adiabatic transition is assumed and the ratio of is changed at high Mikheyev–Smirnov–Wolfenstein resonance and depends on the mass ordering of neutrinos (Smirnov, 2003). Different transition probabilities are assumed for normal and inverted neutrino mass orderings to account for the change in ratio of electron flavour neutrinos due to the dense stellar medium and the effects of neutrino oscillations in vacuum.
@@
We attempt to explore the sensitivities of detecting pre-SN neutrinos from the well-known pre-SN candidate -Ori (Betelgeuse). Current estimation of its mass and distance suggests 16.5-19 M⊙ and pc (Joyce et al., 2020). Limited by the available data sets, we choose 15 M⊙ and 150 pc to simulate a Betelgeuse-like pre-SN star in this work. Although we focus on detecting pre-SN neutrinos from a Betelgeuse-like star, pre-SN neutrinos from other stars could also be observable. A list of candidate pre-SN stars with updated distance and mass estimates can be found in (Machado et al., 2022). Figure 1 shows the expected number of IBD candidates per kton of water for different pre-SN models as a function of the energy, assuming a Betelgeuse-like pre-SN star. The event spectra are obtained by integrating the expected candidates over the last 24 hours, 12 hours, 6 hours and 1 hour prior to core collapse. These candidates predominantly cluster around MeV, but can be found at higher energies. The expected event rate increases over time, leading to a large fraction of the total IBD candidates concentrated in the last hour.
@@ 3 The KamLAND experiment
KamLAND is an LS detector located 1,000 m underground in the Kamioka mine. KamLAND was originally designed to study reactor neutrinos, geoneutrinos, and low-energy solar neutrinos. The primary target volume consists of 1 kton of ultra-pure LS contained in a 13 m diameter spherical balloon made of 135 m-thick transparent nylon ethylene vinyl alcohol copolymer (EVOH) composite film. The components of the KamLAND LS is 80% dodecane and 20% pseudocumene (1,2,4-trimethylbenzene) with 1.36 g/L of the fluor PPO (2,5-diphenyloxazole). An array of 1,325 17-inch photomultiplier tubes (PMTs) and 554 20-inch PMTs mounted on the inner surface of an 18 m diameter stainless steel sphere is used to detect the scintillation light from events occurring within the balloon. Non-scintillating mineral oil fills the space between the balloon and the inner surface of the sphere. This is all surrounded by a 3.2 kton water Cherenkov detector contained in a resin-coated cylindrical rock cavern for cosmic-ray veto. Detailed information of the detector is given in (Suzuki, 2014).
@@
In 2011, an 1.54 m radius inner balloon containing Xe-loaded liquid scintillator (Xe-LS) was installed in the center of the main balloon as a part of the KamLAND Zero-Neutrino Double-Beta Decay (KamLAND-Zen) experiment (Gando et al., 2012a). The inner balloon was updated to have a 1.92 m radius to house double amount of Xe-LS in 2018 (Abe et al., 2023). The center region is not used for the analysis because of backgrounds from the inner balloon and its support materials.
The energy and vertex of an event within the balloon can be reconstructed using the timing and charge distributions of scintillation photons recorded by the PMTs. The reconstruction algorithms are calibrated with on-axis and off-axis radioactive sources deployed from a glove box installed at the top of the detector. Using these calibration sources and naturally occurring radioactive sources, the energy resolution is estimated to be and the vertex resolution is estimated to be , respectively, where is visible energy in the KamLAND detector. The nonlinear and particle-dependent effects of the conversion between deposited (real) energy and are also calibrated with a model incorporating Birks quenching and Cherenkov emission.
@@ 3.1 Event selection in KamLAND
KamLAND performs muon vetos prior to selecting prompt-delayed pairs (delayed coincidence method) as neutrino events. Cosmic-ray muons make bright scintillation events and multiple spallation products, including neutrons. This makes it challenging to reconstruct the correct vertex and energy of low-energy events and to select prompt-delayed pairs immediately following the muon event. Thus, all events within 2 ms of muons are vetoed. However, the 2-ms veto is not enough for showering muons. Showering muons generate non-negligible amount of long-lived spallation products such as 9Li, which has a lifetime of 257.2 ms. Therefore, KamLAND performs a 2-s whole-volume veto for the showering muons (Gando et al., 2012b). Alternatively, a cylindrical cut along the trajectory is applied when the reconstruction quality is good. These three muon vetos are determined by the total observed charge, the residual charge, and the quality of muon event reconstruction.
After applying the muon vetos, KamLAND applies the following criteria: (i) prompt energy: ; (ii) delayed energy: (capture on proton), or (capture on 12C); (iii) spatial correlation between the prompt and delayed events: ; (iv) time difference between prompt and delayed events: ; (v) fiducial volume (FV) radii: ; (vi) inner balloon cut: and for , where is the reconstructed delayed vertex. Note, the prompt energy () is the sum of the positron kinetic energy and annihilation energies.
@@
The total selection efficiency is calculated via the Geant4 MC simulation. events are generated uniformly in the 750 cm radius volume for each prompt energy bin, and the delayed-coincidence selection with the likelihood selection are applied. Then, the selection efficiency is calculated as ratio of the number of survived events after the selection to the number of events generated within the 600 cm FV, and is shown in Figure 2. The efficiency at low energies mirrors the spectrum of the accidental background because of the likelihood-based selection. At high energies, it remains nearly constant. The efficiency loss is dominated by the inner balloon cut. Without this cut, the efficiency is higher, 90 at above 4 MeV. The muon veto has an additional effect to reduce the analysis time. The residual analysis time after this reduction is defined as the livetime. The KamLAND data are divided into runs. The average lifetime ratio in any single run is approximately 0.903.
All PMT waveforms undergo digitization through front-end electronics and are collected by the Data Acquisition (DAQ) software. Event energy and vertex are reconstructed from these waveforms and are collected as a single file, each covering approximately 6 minutes of data. It takes 300–900 s, on average 700 s, from the time of the last event in the file to the end of the reconstruction. The latency is 800–1200 s, on average 1120 s, when we consider the first event in the file. The latency highly depends on other process status. The pre-SN monitoring process is scheduled to run at 5 minute intervals. Upon the identification of a new file, the process applies the selection criteria described above to detect neutrino events. Additionally, the process counts the number of events that passed the selection criteria within the past 24 hours for pre-SN analysis.

### v3

@@ Abstract
Preceding a core-collapse supernova, various processes produce an increasing amount of neutrinos of all flavors characterized by mounting energies from the interior of massive stars. Among them, the electron antineutrinos are potentially detectable by terrestrial neutrino experiments such as KamLAND and Super-Kamiokande via inverse beta decay interactions. Once these pre-supernova neutrinos are observed, an early warning of the upcoming core-collapse supernova can be provided. In light of this, KamLAND and Super-Kamiokande, both located in the Kamioka mine in Japan, have been monitoring pre-supernova neutrinos since 2015 and 2021, respectively. Recently, we performed a joint study between KamLAND and Super-Kamiokande on pre-supernova neutrino detection. A pre-supernova alert system combining the KamLAND detector and the Super-Kamiokande detector was developed and put into operation, which can provide a supernova alert to the astrophysics community. Fully leveraging the complementary properties of these two detectors, the combined alert is expected to resolve a pre-supernova neutrino signal from a 15 M⊙ star within 510 pc of the Earth, at a significance level corresponding to a false alarm rate of no more than 1 per century. For a Betelgeuse-like model with optimistic parameters, it can provide early warnings up to 12 hours in advance.
@@ 1 Introduction
Neutrinos emitted by a supernova during the first 10 seconds carry unique information about the physics of supernovae, which hold immense significance in the realm of astrophysics. The first observed supernova neutrinos (Hirata et al., 1987; Bionta et al., 1987; Alekseev et al., 1987) were from SN1987A in the Large Magellanic Cloud, 50 kpc away from Earth (Pietrzyński et al., 2019). Since then, various neutrino experiments, such as Borexino (Alimonti et al., 2009), IceCube (Köpke, 2018), KamLAND (Abe et al., 2022b), LVD (Agafonova et al., 2015), NOvA (Acero et al., 2020), SNO+ (Andringa et al., 2016), and Super-Kamiokande (SK) (Abe et al., 2016), equipped with advanced technology and improved capabilities, continued the quest to detect supernova neutrino bursts. Furthermore, a number of next-generation neutrino detectors sensitive to galactic supernova neutrinos are under construction, including DUNE (Abi et al., 2021), Hyper-Kamiokande (Abe et al., 2018), JUNO (Abusleme et al., 2023), and KM3NeT (Aiello et al., 2021). To catch such a fleeting event, it is desirable to be alerted well before the explosion, so that astronomers and particle physicists may prepare for the observation of supernova neutrinos and possible gravitational waves as soon as the explosion happens.
The evolution of a single star, whose initial mass is greater than 8 solar masses (M⊙), to its final stages prior to the core-collapse supernova (CCSN) is characterized by nuclear burning in its core due to its high temperature and density (Woosley et al., 2002).111With different physical assumptions, such as rotation, or the presence of a massive companion, the evolution can be significantly different (Eldridge & Stanway, 2022). The change of the chemical composition of a star, forming concentric shells of heavier elements along its volume, is the result of nuclear fusion of heavier elements in the core. Stars at this stage are called pre-supernova (pre-SN) stars. The main cooling mechanism of a pre-SN star is through neutrino emissions. Starting from the carbon burning stage, neutrinos are produced by pair annihilation , producing all flavours of neutrino and antineutrino pairs (Odrzywolek et al., 2004a). As the star approaches core collapse, the nuclear beta decay begins to dominate. Nuclear processes, such as beta decay, will eventually contribute more to the neutrino emission than thermal processes as the star approaches core collapse (Patton et al., 2017).
@@
The energy of pre-SN neutrinos is of sub-MeV scale or MeV scale. We focus on inverse beta decay (IBD) , which has a relatively large cross section in liquid scintillator (LS) detectors (e.g. KamLAND (Suzuki, 2014)) and water Cherenkov detectors (e.g. SK (Fukuda et al., 2003)) in the energy range of pre-SN neutrinos. Compared to LS detectors, whose energy threshold is typically less than 1 MeV, water Cherenkov detectors are less sensitive to low-energy neutrinos because the higher energy threshold, which is 2.5 MeV in kinetic energy as shown in Section 4.1, limits the detection of neutron capture signals. However, starting in 2020, the SK detector was loaded with gadolinium (Gd) to improve the neutron detection efficiency (Abe et al., 2022a). Since 2015, KamLAND has been monitoring pre-SN neutrinos and was able to provide pre-SN alerts to the astrophysics community (Asakura et al., 2016). Later in 2021, SK has also implemented an online pre-SN alert system (Machado et al., 2022). As of May 2024, no alert has been issued. Besides IBD in LS detectors and water Cherenkov detectors, distinct detection methods in other detectors could be utilized for pre-SN neutrino detection as well. For example, coherent neutrino-nucleus scattering in future large scale dark matter direct detection experiments, is a method complementary to IBD, because it can detect all flavors of neutrinos (Raj et al., 2020).
In this article, we introduce a joint study between KamLAND and SK on pre-SN neutrino detection. This combination aims at extending the reach to potential CCSN progenitors at a greater distance and reducing the warning time of pre-SN alerts. Additionally, we present new sensitivity results individually for KamLAND and SK. Compared to the previous study (Asakura et al., 2016), KamLAND has now taken more recent pre-SN neutrino models into consideration. The sensitivity of SK to pre-SN neutrinos with Gd by mass is presented in (Machado et al., 2022). Since 2022, the Gd concentration in SK has increased to 0.03 by mass, further enhancing its capability to identify low-energy electron antineutrinos () via IBD. Therefore, both experiments have reoptimized their analysis strategies according to these changes.
@@ 2 Pre-supernova neutrino model
To estimate the expected signal from pre-SN in SK and KamLAND, two models for pre-SN neutrino emission during the evolution of massive stars were used: (Odrzywolek & Heger, 2010; Odrzywolek et al., 2004b) and (Patton et al., 2017). Both models provide data sets for the calculation of emission during the pre-SN stage. (Odrzywolek et al., 2004b) provides data sets for stars with 15 M⊙ and 25 M⊙ and (Patton et al., 2017) for 15 M⊙ and 30 M⊙.
@@
The model from (Odrzywolek & Heger, 2010) assumes that the entire neutrino flux comes from pair annihilation. For the nuclear isotopic composition of the star, the model assumes a nuclear statistical equilibrium (NSE), which is a treatment only dependent on the temperature, density, and electron fraction, making it a simple flux estimated by only post-processing an already existing stellar model. The model from (Patton et al., 2017) includes a more complete evaluation of the neutrino flux from the pre-SN star, including contributions not only from pair annihilation, but also from plasmon decay, photoneutrino process, beta decay, and electron/positron captures. By using the star evolution code MESA (Modules for Experiments in Stellar Astrophysics) (Paxton et al., 2011), this model couples the isotopic evolution to the stellar evolution, giving a more robust estimation of the neutrino fluxes from nuclear weak processes.
@@
Neutrinos undergo flavor conversion, i.e. neutrino oscillations, from the point of production to the point of detection. To calculate the expected signal from the considered models, adiabatic neutrino oscillations in the matter of the star, and neutrino oscillations in vacuum are taken into account. For the former, the ratio of is changed at high Mikheyev–Smirnov–Wolfenstein resonance, which depends on the mass ordering of neutrinos (Smirnov, 2003). Different transition probabilities are assumed for normal and inverted neutrino mass orderings to account for the change in ratio of electron flavour neutrinos due to the dense stellar medium and the effects of neutrino oscillations in vacuum.
@@
We attempt to explore the sensitivities for detecting pre-SN neutrinos from the well-known red supergiant -Ori (Betelgeuse), which will potentially end up a CCSN.222There are claims that the explosion is imminent (Saio et al., 2023) and opposing views (Molnár et al., 2023). Current estimation of its mass and distance suggests 16.5-19 M⊙ and pc (Joyce et al., 2020). Limited by the available data sets, we choose 15 M⊙ and 150 pc to simulate a Betelgeuse-like pre-SN star in this work. Although we focus on detecting pre-SN neutrinos from a Betelgeuse-like star, pre-SN neutrinos from other stars could also be observable. A list of candidate pre-SN stars with updated distance and mass estimates can be found in (Machado et al., 2022). Figure 1 shows the expected number of IBD candidates per kton of water for different pre-SN models as a function of the energy, assuming a Betelgeuse-like pre-SN star. The event spectra are obtained by integrating the expected candidates over the last 24 hours, 12 hours, 6 hours and 1 hour prior to core collapse. These candidates predominantly cluster around MeV, but can be found at higher energies. The expected event rate increases over time, leading to a large fraction of the total IBD candidates concentrated in the last hour.
@@ 3 The KamLAND experiment
KamLAND is an LS detector located 1,000 m underground in the Kamioka mine. KamLAND was originally designed to study reactor neutrinos, geoneutrinos, and low-energy solar neutrinos. The primary target volume consists of 1 kton of ultra-pure LS contained in a 13 m diameter spherical balloon made of 135 m-thick transparent nylon ethylene vinyl alcohol copolymer (EVOH) composite film. The components of the KamLAND LS are 80% dodecane and 20% pseudocumene (1,2,4-trimethylbenzene) with 1.36 g/L of the fluor PPO (2,5-diphenyloxazole). An array of 1,325 17-inch photomultiplier tubes (PMTs) and 554 20-inch PMTs mounted on the inner surface of an 18 m diameter stainless steel sphere is used to detect the scintillation light from events occurring within the balloon. Non-scintillating mineral oil fills the space between the balloon and the inner surface of the sphere. This is all surrounded by a 3.2 kton water Cherenkov detector contained in a resin-coated cylindrical rock cavern for cosmic-ray veto. Detailed information of the detector is given in (Suzuki, 2014).
@@
In 2011, an inner balloon of 1.54 m radius containing Xe-loaded liquid scintillator (Xe-LS) was installed in the center of the main balloon as a part of the KamLAND Zero-Neutrino Double-Beta Decay (KamLAND-Zen) experiment (Gando et al., 2012a). The inner balloon was updated to have a 1.92 m radius to house double the amount of Xe-LS in 2018 (Abe et al., 2023b). The center region is not used for the analysis because of backgrounds from the inner balloon and its support materials.
The energy and vertex of an event can be reconstructed using the timing and charge distributions of scintillation photons recorded by the PMTs. The reconstruction algorithms are calibrated with radioactive sources deployed from the top of the detector (Berger et al., 2009; Banks et al., 2015). Using these calibration sources, the energy resolution is estimated to be and the vertex resolution is estimated to be , respectively, where is reconstructed energy. The nonlinear and particle-dependent effects of the conversion between deposited (real) energy and are also calibrated with the Birk’s formula (Birks, 1951) and contribution of Cherenkov emission.
@@ 3.1 Event selection in KamLAND
KamLAND performs muon vetos prior to selecting prompt-delayed pairs (delayed coincidence method) as neutrino events. Cosmic-ray muons produce events with bright scintillation light and multiple spallation products, including neutrons. This makes it challenging to reconstruct the correct vertex and energy of low-energy events and to select prompt-delayed pairs immediately following the muon event. Thus, all events within 2 ms of the arrival time of muons are vetoed. However, the 2-ms veto is not enough for high energy muons which make cascade showers in the detector. Such muons generate a non-negligible amount of long-lived spallation products such as 9Li, which has a lifetime of 257.2 ms. Therefore, KamLAND performs a 2-s whole-volume veto for high energy muons (Gando et al., 2012b). Alternatively, a cylindrical cut along the trajectory is applied when the reconstruction quality is good. These three muon vetos are determined by the total observed charge, the residual charge, which means the difference between the observed charge minus the charge that would be expected if the muon simply penetrated the detector, and the quality of muon event reconstruction.
After applying the muon vetos, KamLAND applies the following criteria: (i) reconstructed prompt energy: ; (ii) reconstructed delayed energy: (capture on proton), or (capture on 12C); (iii) spatial correlation between the prompt and delayed events: ; (iv) time difference between prompt and delayed events: ; (v) fiducial volume (FV) radii: ; (vi) inner balloon cut: and for , where is the reconstructed delayed vertex. Note, the reconstructed prompt energy () is the sum of the positron kinetic energy and annihilation energies with the quenching effect.
@@
The total selection efficiency is calculated using a Geant4 Monte-Carlo simulation. A total of events are generated uniformly in the 750 cm radius volume for each prompt energy bin, and the delayed-coincidence selection with the likelihood selection are applied. The selection efficiency is calculated as the ratio of the number of surviving events after the selection to the number of events generated within the 600 cm FV, shown in Figure 2. The efficiency at low energies mirrors the spectrum of the accidental background because of the likelihood-based selection. At high energies, it remains nearly constant. The efficiency loss is dominated by the inner balloon cut. Without this cut, the efficiency is higher, 90 above 4 MeV. The muon veto has an additional effect of reducing the analysis time. The residual analysis time after this reduction is defined as the livetime. The KamLAND data are divided into runs. The average lifetime ratio in any single run is approximately 0.903.
All PMT waveforms undergo digitization through front-end electronics and are collected by the Data Acquisition (DAQ) software. Event energy and vertex are reconstructed from these waveforms and are collected as a single file, each covering approximately 6 minutes of data. It takes 300–900 s, on average 700 s, from the time of the last event in the file to the end of the reconstruction. The latency is 800–1200 s, on average 1120 s, when we consider the first event in the file. The latency strongly depends on the status of other processes. The pre-SN monitoring process is scheduled to run at 5 minute intervals. Upon the identification of a new file, the process applies the selection criteria described above to detect neutrino events. Additionally, the process counts the number of events that passed the selection criteria within the past 24 hours for pre-SN analysis.